OpenStudy (anonymous):

in how many ways can 6 people be seated in a motor car if only 2 can occupy the driver's position?

OpenStudy (anonymous):

Do you remember what a permutation is?

OpenStudy (anonymous):

i believe so, but i wouldn't mind if you reminded me of what it was hehehe

OpenStudy (anonymous):

Well a permutation is the number of ways a group of objects can be arranged.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

and order matters

OpenStudy (earthcitizen):

6P2

OpenStudy (anonymous):

@ earth citizen, are you sure because my solution booklet says the answer is 240

OpenStudy (anonymous):

but i am not sure how they got that, does anyone know and if you do could you please show worked solutions, thank you!

OpenStudy (anonymous):

Sorry electrict failure.

OpenStudy (anonymous):

I see this question is very popular. So i'll take a seat.

OpenStudy (anonymous):

haha nice pun there

OpenStudy (earthcitizen):

is it repeated or non-repeated ?

OpenStudy (anonymous):

it doesn't specify

OpenStudy (anonymous):

\[\left(\begin{matrix}2\\ 1\end{matrix}\right)\left(\begin{matrix}5 \\ 5\end{matrix}\right)\]

OpenStudy (turingtest):

For every time you have one of the two given people in the driver's seat the number of permutations of the remaining seats is 5! there are two possible drivers, thus the total number of permutations is 5!+5!=240

OpenStudy (anonymous):

roachie i don't understand what you have written, would it be ok if you explained thanks

OpenStudy (anonymous):

6 riders 2 can drive , at any time there are 5 passengers, 5 combinations of five is 120 , 1 combination of 2 is 3 120 * 2 = 240

OpenStudy (anonymous):

oh i see thank you turing test!!

OpenStudy (turingtest):

very welome

OpenStudy (turingtest):

welcome* lol

OpenStudy (anonymous):

ahhh thank you for the clarification roachie! and Lol @ turing test's typo hehhehe